OpenStudy (vijeya3):

The moon has mass M and radius R. A small object is dropped from a distance of 3R from the moon’s center. The object’s impact speed when it strikes the surface of the moon is equal to (kGM/R)^(1/2) for k =

imqwerty (imqwerty):

find the gravity's equation and then use this-> mgh=1/2 mv^2 at top position it will have only PE and when it will come to surface all the PE will convert into KE its not alright tho..because there is a certain velocity after which the velocity will stop increasing(terminal velocity) <-we don't consider this because because no drag forces are specified + the gravity is not same everywhere but we assume it to be constant

Parth (parthkohli):

really? I don't recall them teaching us this form of potential energy in class 9. weird.

imqwerty (imqwerty):

they taught us tho you talking about mgh ?

Parth (parthkohli):

no, they taught that but I think it's better to use -GMm/R here

imqwerty (imqwerty):

oh okay :)

Parth (parthkohli):

yeah I wonder where she got this problem

OpenStudy (vijeya3):

@ParthKohli I got this problem from here- http://www.cracksat.net/sat2/physics/test603.html

OpenStudy (vijeya3):

And yeah/...9th problems are damn simple and boring...

imqwerty (imqwerty):

take it to be 3R because we consider distance from the center

imqwerty (imqwerty):

and well when it will strike the ground we must also consider mgR as PE 

Parth (parthkohli):

well then you'll need to know either of these:\[g=-\frac{GM}{R^2}\]\[\text{GPE} = \frac{-GMm}{R}\]

OpenStudy (vijeya3):

yeah..i took 3R and i am getting the wrong answer...please ek baar tum calculate karo na aayush...

OpenStudy (vijeya3):

yeah parth..i did that...

OpenStudy (vijeya3):

sorry i got 4/9* and correct is 4/3(that is what the site says)

imqwerty (imqwerty):

\(g=\frac{-GM}{R^2}\) \(mg(3R)=\frac{1}{2}mv^2 + mgR\) \(mg(2R)= \frac{1}{2}mv^2\) \( \frac{-GM}{R^2} (2R)= \frac{1}{2} v^2\) \(\left(\frac{-4GM}{R} \right)^{\frac{1}{2}}=v\)

imqwerty (imqwerty):

4/3 :/

Parth (parthkohli):

\[\underbrace{\frac{1}2 mv^2}_{\text{increase in kinetic energy}} = \underbrace{GMm\left(\frac{1}R - \frac{1}{3R}\right)}_{\text{decrease in potential energy}}\]

OpenStudy (vijeya3):

okay..I was taking 3R in g ka formula and 2R for h in mgh...I guess that is where I erred...

OpenStudy (vijeya3):

But aayush...while finding g,don't we take center to center distance?

imqwerty (imqwerty):

okay we had to take different gravities \(g_{at~top}=\frac{-GM}{(3R)^2}\) \(g_{at~surface}=\frac{-GM}{(R)^2}\) \(mg_{top}(3R)=\frac{1}{2}mv^2 + mg_{surface} R\) \(m \frac{-GM}{(3R)^2}(3R)-m\frac{-GM}{(R)^2}R= \frac{1}{2}mv^2\) coming to same equations as pk's

Parth (parthkohli):

stfu

imqwerty (imqwerty):

yes

OpenStudy (vijeya3):

different gravities?